# Performance Assessment of Universal Machine Learning Interatomic Potentials: Challenges and Directions for Materials’ Surfaces

### v1

@@ Abstract
Machine learning interatomic potentials (MLIPs) are one of the main techniques in the materials science toolbox, able to bridge ab initio accuracy with the computational efficiency of classical force fields. This allows simulations ranging from atoms, molecules, and biosystems, to solid and bulk materials, surfaces, nanomaterials, and their interfaces and complex interactions. A recent class of advanced MLIPs, which use equivariant representations and deep graph neural networks, is known as universal models. These models are proposed as foundational models suitable for any system, covering most elements from the periodic table. Current universal MLIPs (UIPs) have been trained with the largest consistent dataset available nowadays. However, these are composed mostly of bulk materials’ DFT calculations. In this article, we assess the universality of all openly available UIPs, namely MACE, CHGNet, and M3GNet, in a representative task of generalization: calculation of surface energies. We find that the out-of-the-box foundational models have significant shortcomings in this task, with errors correlated to the total energy of surface simulations, having an out-of-domain distance from the training dataset. Our results show that while UIPs are an efficient starting point for fine-tuning specialized models, we envision the potential of increasing the coverage of the materials space towards universal training datasets for MLIPs.
@@ I Introduction
Recently, a class of MLIPs employing deep graph networks [12, 13, 14] has improved previous results in a range of systems and benchmarks. With the use of E(3)-equivariant representations, models have achieved not only accuracy but also unprecedented data efficiency, such as the Neural Equivariant Interatomic Potentials (NequIP) [15]. These advances led the community to envision universal MLIPs (UIPs), general models able to simulate any system across most of the periodic table [16]. These universal potentials have been proposed as foundational models, a starting point from which further specialized models could be trained upon.
Chronologically, these models improved by growing both their datasets and number of parameters: M3GNet [17] was trained with 188k structures and has 228k params; CHGNet [18] uses atomic magnetic moments as additional representation inputs and introduced the MPtrj dataset with 1.58M structures from the Materials Project database and has 413k params; MACE MP-0 [19, 20, 21] was also trained with MPtrj and has 4.69M params; and the proprietary GNoME model [22] was trained with an active learning dataset of 89M structures starting from the Materials Project, and has 16.2M parameters. As usual in the deep learning community, a catalyst for advances is the creation of benchmarks to assess performance in a comparable task, the Matbench Discovery [23] serves this purpose in the task of materials discovery for universal models. However, the models are trained on the largest systematic DFT databases available to date, which include millions of bulk materials simulations but are currently not comprehensible in the complete materials space of all possible combinations of atoms, compositions, and structures, such as molecules, surfaces and interfaces, nanomaterials, and so on.
@@ II.1 Dataset exploration
We start by querying the Materials Project (MP) database for data on surfaces. This data is available for unary systems [24]. We gathered a total of 1497 different surface structures that are generated from a total of 138 different bulk systems, comprehending 73 different chemical elements.
@@ II.2 Inference of universal machine learning atomistic potentials
We proceed with an assessment of the universal MLIPs (UIPs) in their ability to predict the surface energy data of the materials under consideration. Initially, we gauge their performance in predicting the total energy of the bulk structures. Given their nature as universal MLIPs trained on bulk data sourced from the Materials Project, a high level of accuracy is anticipated for these specific structures.
[…]
Figure 2b illustrates the prediction error for the total energy of surfaces. As these structures differ from the models’ training set, the error rate increases. Surprisingly, CHGNet exhibits greater accuracy than both MACE and M3GNet in this context, with M3GNet also surpassing MACE. This observation highlights the inherent trade-off between accuracy and generalization. Although MACE performs better on the test set, its predictive capability diminishes when extrapolating to unknown data dissimilar to the training set.
Figure 2: Performance assessment of the universal interatomic potentials over the surfaces dataset. (a) Parity plot for the total energy per atom of the bulk systems that gave origin to the surfaces of the dataset. (b) Parity plot for the total energy per atom of the surfaces within the dataset. (c) Parity plot for the surface energy ($\gamma_{hkl}^{\sigma}$) for the surfaces within the dataset. For (a), (b), and (c), the dashed line marks the $x=y$ line. (d) Boxplot and violin plot for the error in the prediction of the surface energy ($\gamma_{hkl}^{\sigma}$) from the three universal interatomic potentials evaluated. The horizontal lines in the middle of the boxes mark the medians. The boxes are plotted from the first to the third quartile. The whiskers extend to $1.5$ times the interquartile range from the third (above) and first (below) quartile and empty circles mark the outliers (data points outside the whiskers).
@@
Figure 2c shows the parity plot for these models and Fig. 2d their error distribution. In the case of surface energy, CHGNet is still the most accurate, however MACE follows in second and M3GNet in third. CHGNet mostly underestimates the surface energy while MACE shows several cases of overestimation. Like CHGNet, M3GNet underestimates the surface energy. Ideally, accuracy over bulk and surface total energy is desired to achieve accurate surface energies, however, error cancellation also plays a part in this analysis, therefore keeping in comparison the bulk and surface total energy is essential to gain insight into these models’ performance. Additionally, we do not observe any correlation between the predictions of each model (see the Supplemental Information), that is, the errors are inherent from each model’s limits due to their different training set, representations, and architecture.
@@ II.3.1 System chemistry
Next, we show the RMSE of these models separated by system chemistry. Figure 3 shows a comparison between the RMSE on the surface energy and also bulk and surface total energy per atom, which we take as the limiting cases for the accuracy of each model. Figure 3 illustrates that errors in bulk predictions are not consistently correlated with errors in surface energy predictions. For example, while MACE shows poor performance in predicting the surface energy of tungsten (W), its bulk predictions are relatively accurate. Conversely, for caesium (Cs), both bulk and surface total energy predictions are accurate, but the surface energy predictions are not, which is a result of error cancellation. Despite these discrepancies, MACE demonstrates small errors for several other systems, for instance, copper (Cu) is well represented by MACE. Likewise, M3GNet exhibits a higher frequency of large errors, particularly for elements such as boron (B), carbon (C), manganese (Mn), iron (Fe), niobium (Nb), molybdenum (Mo), technetium (Tc), ruthenium (Ru), rhenium (Re), osmium (Os), iridium (Ir), and platinum (Pt). CHGNet shows fewer problematic predictions compared to M3GNet. Moreover, elements such as carbon (C), manganese (Mn), tungsten (W), iron (Fe), rhenium (Re), osmium (Os), and iridium (Ir) also require careful consideration, as these elements also present challenges for the UIPs. The comparison with the surface total energy shows that they are the major root for the deviations of the surface energy predictions, as expected. Surfaces are not included in the training set of these universal models, therefore they struggle with extrapolating from the bulk structures.
@@ II.3.3 System local geometries (local environments, beyond surfaces structures)
Figure 5 shows the kPCA map colored by the absolute error given by the MACE model for surface energy. Interestingly, the MACE representation is enough to single out most systems. The representation seems to take great advantage of the different chemical species within the dataset. As shown in Fig. 3, tungsten (W) presents most of the error from MACE. Additionally, in Fig. 5 we also try to associate the kPCA map with the lattice of each original bulk structure. We highlight the surfaces that originate from face-centered cubic (FCC), body-centered cubic (BCC), and hexagonal closed-packed (HCP) lattices. The error is not directly dependent on the lattice symmetry. Detailed analysis of each system chemistry on the kPCA map is available in the Supporting Information.
@@ II.4.1 Universal vs NequIP/MTP
Our next step in this assessment is to address the question: “If one is interested in materials’ surfaces, are universal interatomic potentials a good starting point, or should one train a specialized interatomic potential for the task at hand?”. For this comparison, we started by selecting candidates for specialized MLIPs. Keeping in mind the performance-computational cost trade-off, the moment tensor potential (MTP) [27] formalism is ideal since it can attain very low test errors with a rather small dataset and number of parameters [25, 28]. Zuo et al. [25] performed a systematic benchmark of “pre-universal” MLIPs and made available not only the training and test sets of their performance and computational cost assessment, but also the models they have trained for the elements: nickel (Ni), copper (Cu), lithium (Li), molybdenum (Mo), silicon (Si), and germanium (Ge). We called MTP [JPCA] the set of MTP models as available from ref. [25] for each of the elements considered in their work. We trained another set of MTP models for these elements, however, using the bulk structures as queried from the Materials Project database. We called this set of models MTP [$*{}^{*}$], as this is derived from a reduced version of the MPtrj dataset, that is, for these selected elements, as in ref. [25]. Mind that for both MTP [JPCA] and MTP [$*{}^{*}$], there is a model for each element, while for UIPs, a single model handles all the elements.
@@
We added the Neural Equivariant Interatomic Potential (NequIP) to our comparison as they are equivariant neural networks, similar to the UIPs, and highly efficient in data use. In this case, we trained two sets of NequIP models, one using ref. [25] training and test sets, which we called NequIP [JPCA], and a second set of models, as we did with MTP, using the bulk structures as queried from the Materials Project for the same elements used in ref. [25], which we call NequIP [$*{}^{*}$]. Our final comparison contrasts the UIPs: MACE, M3GNet, and CHGNet with four sets of specialized MLIPs: MTP [JPCA], MTP [$*{}^{*}$], NequIP [JPCA], and NequIP [$*{}^{*}$], as shown in Fig. 6 and summarized in Table 1. See the Supporting Information for the parity plot containing all these models for this reduced set of elements.
@@
First, by comparing the RMSE in Table 1 with the RMSE of the full set of elements from the MPtrj dataset, the RMSE of UIPs shows a different trend. For this smaller set of elements, although MACE is still not the most accurate total energy prediction on the bulk structures, there is not a significant increase in the error for surface total energy prediction. M3GNet on the other hand shows a significant error increase. CHGNet however follows the opposite trend, with surface total energy errors being smaller than bulk total energy errors. For surface energy, MACE is the most accurate between the UIPs.
@@
Comparing first MTP and NequIP with the two different datasets: JPCA and $*{}^{*}$. The models trained with the JPCA dataset present an overall better performance. We attribute this to be sheer influence of the construction of the dataset. The JPCA dataset is more diverse in types of structures containing: MD trajectory snapshots, defects, deformations, and even a few surfaces, while the $*{}^{*}$ datasets contain only bulk structures at their relaxed geometries and relaxation trajectories. Comparison between MTP and NequIP for the same dataset brings the evident power of MTP. Although NequIP showcases smaller RMSE, MTP is not far behind, especially for bulk structures. However, as we go to surfaces, NequIP showcases higher generalization power. Comparison between specialized and universal IPs demonstrates that UIPs lack the generalization power to yield reasonable predictions for surfaces. This is the case when taking into consideration both MTP and NequIP models trained with the JPCA dataset. As expected, the lack of diversity on the $*{}^{*}$ dataset hinders the performance of specialized MLIPs.
@@
Figure 6: Performance comparison between universal, specialized, and universal fine-tuned interatomic potentials. The root mean squared errors (RMSE) of the universal interatomic potentials are compared in relation to the surface chemistry, for selected elements. The upper triangle is the surface energy ($\gamma_{hkl}^{\sigma}$) RMSE, the lower leftmost triangle is the bulk total energy per atom RMSE and the lower rightmost triangle is the surface total energy per atom RMSE. The numbers in parenthesis are the number of surface structures evaluated for each chemical element.
@@
Table 1: Root mean squared errors (RMSE) for universal interatomic potentials, specialized machine learning interatomic potentials, and fine-tuned universal interatomic potential for a selected set of elements: Ni, Cu, Li, Mo, Si, and Ge. For the column type, we label the models as universal interatomic potential (UIP), local representation interatomic potential (LRIP), graph neural network (GNN), and UIP fine-tuned (UIP-FT).

### v2

@@ Abstract
Machine learning interatomic potentials (MLIPs) are one of the main techniques in the materials science toolbox, able to bridge ab initio accuracy with the computational efficiency of classical force fields. This allows simulations ranging from atoms, molecules, and biosystems, to solid and bulk materials, surfaces, nanomaterials, and their interfaces and complex interactions. A recent class of advanced MLIPs, which use equivariant representations and deep graph neural networks, is known as universal models. These models are proposed as foundation models suitable for any system, covering most elements from the periodic table. Current universal MLIPs (UIPs) have been trained with the largest consistent dataset available nowadays. However, these are composed mostly of bulk materials’ DFT calculations. In this article, we assess the universality of all openly available UIPs, namely MACE, CHGNet, and M3GNet, in a representative task of generalization: calculation of surface energies. We find that the out-of-the-box foundation models have significant shortcomings in this task, with errors correlated to the total energy of surface simulations, having an out-of-domain distance from the training dataset. Our results show that while UIPs are an efficient starting point for fine-tuning specialized models, we envision the potential of increasing the coverage of the materials space towards universal training datasets for MLIPs.
@@ I Introduction
Recently, a class of MLIPs employing deep graph networks [12, 13, 14, 15] has improved previous results in a range of systems and benchmarks. With the use of E(3)-equivariant representations, models have achieved not only accuracy but also unprecedented data efficiency, such as the Neural Equivariant Interatomic Potentials (NequIP) [16]. These advances led the community to envision universal MLIPs (UIPs), general models able to simulate any system across most of the periodic table [17]. These universal potentials have been proposed as foundational models, a starting point from which further specialized models could be trained upon.
Chronologically, these models improved by growing both their datasets and number of parameters: M3GNet [18] was trained with 188k structures and has 228k params; CHGNet [19] uses atomic magnetic moments as additional representation inputs and introduced the MPtrj dataset with 1.58M structures from the Materials Project database and has 413k params; MACE MP-0 [20, 21, 22] was also trained with MPtrj and has 4.69M params; the proprietary GNoME model [23] was trained with an active learning dataset of 89M structures starting from the Materials Project, and has 16.2M parameters; and the recent proprietary MatterSim models [24] use a dataset of 17M structures and can reach up to 182M parameters. As usual in the deep learning community, a catalyst for advances is the creation of benchmarks to assess performance in a comparable task, the Matbench Discovery [25] serves this purpose in the task of materials discovery for universal models. However, the models are trained on the largest systematic DFT databases available to date, which include millions of bulk materials simulations but are currently not comprehensible in the complete materials space of all possible combinations of atoms, compositions, and structures, such as molecules, surfaces and interfaces, nanomaterials, and so on.
@@ II.1 Dataset exploration
We start by querying the Materials Project (MP) database for data on surfaces. This data is available for unary systems [26]. We gathered 1497 different surface structures that were generated from 138 different bulk systems, comprehending 73 different chemical elements.
@@ II.2 Inference of universal machine learning atomistic potentials
We proceed with an assessment of the universal MLIPs (UIPs) in their ability to predict the surface energy data of the materials under consideration. In this work we focus on using the following UIPs: MACE [22], CHGNet [19], and M3GNET [18]. For MACE we use the latest $L=2$ version, for CHGNet we use the 0.3.0 version, and for M3GNet we use the updated version trained with DIRECT sampling of the dataset [27]. Specific version details can be found in Table 2.
Initially, we gauge their performance in predicting the total energy of the bulk structures. Given their nature as universal MLIPs trained on bulk data sourced from the Materials Project, a high level of accuracy is anticipated for these specific structures. Figure  2a shows the performance of UIPs on the bulk systems that give origin to the surfaces of the dataset. As anticipated, the universal models demonstrate proficiency in predicting the total energy of bulk structures. However, their performance falls short of the precision achievable by specialized models tailored to individual systems, which in turn are expected to be below 1 meV atom-1 for selected systems [28]. In contrast, the Matbench Discovery [25] indicates that these universal models yield RMSE values close to 100 meV atom-1 on their test data.
Figure 2b illustrates the prediction error for the total energy of surfaces. As these structures differ from the models’ training set, the error rate increases. Surprisingly, CHGNet exhibits greater accuracy than both MACE and M3GNet in this context, with MACE surpassing M3GNet. This observation highlights the inherent trade-off between accuracy and generalization as M3GNet’s predictive capability diminishes when extrapolating to unknown data dissimilar to the training set.
Figure 2: Performance assessment of the universal interatomic potentials over the surfaces dataset. (a) Parity plot for the total energy per atom of the bulk systems that gave origin to the surfaces of the dataset. (b) Parity plot for the total energy per atom of the surfaces within the dataset. (c) Parity plot for the surface energy ($\gamma_{hkl}^{\sigma}$) for the surfaces within the dataset. For (a), (b), and (c), the dashed line marks the $x=y$ line. (d) Boxplot and violin plot for the error ($\gamma_{\rm ML}-\gamma_{\rm DFT}$) in the prediction of the surface energy ($\gamma_{hkl}^{\sigma}$) from the three universal interatomic potentials evaluated. The horizontal lines in the middle of the boxes mark the medians. The boxes are plotted from the first to the third quartile. The whiskers extend to $1.5$ times the interquartile range from the third (above) and first (below) quartile and empty circles mark the outliers (data points outside the whiskers).
@@
Figure 2c shows the parity plot for these models and Fig. 2d their error distribution. In the case of surface energy, MACE is the most accurate, however CHGNet follows in second and M3GNet in third. The three UIPs mostly underestimate the surface energy with few cases of overestimation. This behavior is reported as a “softening” of the potential energy surface (PES) of UIPs, also reported by Deng et al.[29]. Ideally, accuracy over bulk and surface total energy is desired to achieve accurate surface energies, however, error cancellation also plays a part in this analysis, therefore keeping in comparison the bulk and surface total energy is essential to gain insight into these models’ performance. Additionally, we do not observe any correlation between the predictions of each model (see the Supplemental Information), that is, the errors are inherent from each model’s limits due to their different training set, representations, and architecture.
@@ II.3.1 System chemistry
Next, we show the RMSE of these models separated by system chemistry. Figure 3 shows a comparison between the RMSE on the surface energy and also bulk and surface total energy per atom, which we take as the limiting cases for the accuracy of each model. Figure 3 illustrates that errors in bulk predictions are not consistently correlated with errors in surface energy predictions. For example, while MACE shows poor performance in predicting the surface energy of iron (Fe), its bulk predictions are relatively accurate. MACE demonstrates small errors for several other systems, for instance, copper (Cu) is well represented by MACE on average. Likewise, the three UIPs exhibit a higher frequency of large errors, particularly for elements such as boron (B), carbon (C), manganese (Mn), iron (Fe), niobium (Nb), molybdenum (Mo), technetium (Tc), ruthenium (Ru), rhenium (Re), osmium (Os), iridium (Ir), and platinum (Pt). Both MACE and CHGNet show fewer problematic predictions compared to M3GNet. Moreover, elements such as carbon (C), manganese (Mn), tungsten (W), iron (Fe), rhenium (Re), osmium (Os), and iridium (Ir) also require careful consideration, as these elements also present challenges for the UIPs.
The comparison with the surface total energy shows that they are the major root for the deviations of the surface energy predictions, as expected. Surfaces are not included in the training set of these universal models, therefore they struggle with extrapolating from the bulk structures. Also, since the three models share difficulties in certain elements, these errors are expected to be derived from the dataset. It is also impressive that a much smaller model M3GNet (227.5k parameters) can provide predictions as accurate as MACE and CHGNet (5.7M and 412.5k parameters, respectively) for most elemental systems.
@@ II.3.3 System local geometries (local environments, beyond surfaces structures)
Figure 5 shows the kPCA map colored by the absolute error given by the MACE model for surface energy. Interestingly, the MACE representation is enough to single out most systems. The representation seems to take great advantage of the different chemical species within the dataset. As shown in Fig. 3, osmium (Os) presents most of the error from MACE. Additionally, in Fig. 5 we also try to associate the kPCA map with the lattice of each original bulk structure. We highlight the surfaces that originate from face-centered cubic (FCC), body-centered cubic (BCC), and hexagonal closed-packed (HCP) lattices. The error is not directly dependent on the lattice symmetry. Detailed analysis of each system chemistry on the kPCA map is available in the Supporting Information.
@@ II.4.1 Universal vs NequIP/MTP
Our next step in this assessment is to address the question: “If one is interested in materials’ surfaces, are universal interatomic potentials a good starting point, or should one train a specialized interatomic potential for the task at hand?”. For this comparison, we started by selecting candidates for specialized MLIPs. Keeping in mind the performance-computational cost trade-off, the moment tensor potential (MTP) [31] formalism is ideal since it can attain very low test errors with a rather small dataset and number of parameters [28, 32]. Zuo et al. [28] performed a systematic benchmark of “pre-universal” MLIPs and made available not only the training and test sets of their performance and computational cost assessment, but also the models they have trained for the elements: nickel (Ni), copper (Cu), lithium (Li), molybdenum (Mo), silicon (Si), and germanium (Ge). We called MTP [Ong21] the set of MTP models as available from ref. [28] for each of the elements considered in their work.
@@
We added the Neural Equivariant Interatomic Potential (NequIP) to our comparison as they are equivariant neural networks, similar to the UIPs, and highly efficient in data use. We trained a NequIP model using ref. [28] training and test sets, which we called NequIP [Ong21]. Our comparison contrasts the UIPs: MACE, M3GNet, and CHGNet with two sets of specialized MLIPs: MTP [Ong21], and NequIP [Ong21], as shown in Fig. 6 and summarized in Table 1. See the Supporting Information for the parity plot containing all these models for this reduced set of elements.
@@
First, by comparing the RMSE in Table 1 with the RMSE of the full set of elements from the MPtrj dataset, the RMSE of UIPs shows a different trend. For this smaller set of elements, although MACE and CHGNet are still not the most accurate total energy prediction on the bulk structures, there is not a significant increase in the error for surface total energy predictions, in fact, the RMSE decreases for the prediction of surface total energy. M3GNet on the other hand shows an error increase for the surfaces. For surface energy per area, MACE shows the smallest error among the UIPs.
@@
We will now compare the MTP [Ong21] and NequIP [Ong21] models with UIPs. These models trained with the [Ong21] dataset present an overall better performance. We attribute this to the sheer influence of the construction of the dataset. The [Ong21] dataset is more diverse in types of structures, containing: MD trajectory snapshots, defects, deformations, and even a few surfaces, while the MPtrj, although very large datasets contains only bulk structures close to their relaxed geometries and relaxation trajectories. Comparison between MTP and NequIP brings the evident power of MTP. Although NequIP showcases smaller RMSE, MTP is not far behind, especially for bulk structures. However, as we go to surfaces, NequIP showcases higher generalization power. Comparison between specialized and universal IPs demonstrates that UIPs could have their generalization power improved to yield predictions for surfaces with sufficient accuracy.
@@
Figure 6: Performance comparison between universal, specialized, and universal fine-tuned interatomic potentials. The root mean squared errors (RMSE) of the universal interatomic potentials are compared to the surface chemistry, for selected elements. The upper triangle is the surface energy ($\gamma_{hkl}^{\sigma}$) RMSE, the lower leftmost triangle is the bulk total energy per atom RMSE and the lower rightmost triangle is the surface total energy per atom RMSE. The numbers in parenthesis are the number of surface structures evaluated for each chemical element.
@@
Table 1: Root mean squared errors (RMSE) for universal interatomic potentials, specialized machine learning interatomic potentials, and fine-tuned universal interatomic potential for a selected set of elements: Ni, Cu, Li, Mo, Si, and Ge. For the column type, we label the models as universal interatomic potential (UIP), UIP fine-tuned (UIP-FT), graph neural network (GNN), and moment tensor potentials (MTP).
@@
(eV atom-1)
Surface total energy
(eV atom-1)
Surface energy
(J m-2)
Type
MACE 0.0856 0.0455 0.3512 UIP
CHGNet 0.0791 0.0386 0.5128 UIP
M3GNet 0.0084 0.0704 0.4246 UIP
MACE [FT] 0.0464 0.0139 0.2228 UIP-FT
M3GNet [FT] 0.0399 0.0590 0.3852 UIP-FT
MACE [Ong21] 0.0372 0.0580 0.2144 GNN
M3GNet [Ong21] 0.0448 0.0859 0.7247 GNN
NequIP [Ong21] 0.0378 0.0542 0.2696 GNN
MTP [Ong21] 0.0371 0.0537 0.2820 MTP
Figure 7: Performance comparison between universal, specialized, and universal fine-tuned interatomic potentials for prediction of relative surface energy order. The evaluated surfaces are derived from the bulk structures mp-23, mp-30, mp-51, mp-129, mp-149, and mp-32, for Ni, Cu, Li, Mo, Si, and Ge, respectively.
As demonstrated in Fig. 6, for most elements where UIPs show an increased error, MTP [Ong21] and NequIP [Ong21] demonstrate high-accuracy predictions. Predictions for systems such as Mo and Si show great improvement from the specialized MLIPs. Systems containing Ge on the other hand show worse performance for bulk structures while slightly improved performance for surface structures, as expected from the results from Zuo et al. [28].
To understand the relevance of the dataset when comparing results between MTP and NequIP trained with [Ong21] dataset, we create another comparison, as detailed in Supporting Information, with the so-called MPtrj. The MPtrj dataset is created by filtering the MPtrj dataset for unary structures containing the following elements: Ni (96 structures), Cu (55 structures), Li (83 structures), Mo (25 structures), Si (559 structures), and Ge (225 structures). A single MTP and NequIP model is fitted for each of these elements and compared to the [Ong21] dataset. As expected, since this new dataset has a much less diverse set of structures, with forces and stresses close to zero, these non-universal models struggle to predict accurate surface (total) energies, particularly MTP. The full comparison with the MPtrj dataset and further details can be found in the Supporting Information.
II.4.2 Universal (zero-shot) vs Universal (fine-tuned) vs From scratch models
Since UIPs seem to not be great at dealing with surfaces, we now address the question: “What can we do to improve them?”. One route to improving general models such as the UIPs is fine-tuning, which consists of continuing the training of these models however with a different dataset, tailored to make these universal models more specialized in particular downstream tasks. In our case, we can use the training and test sets of ref. [28] to fine-tune these UIPs. By introducing a more diverse set of structures for each chemical element, we expect that the UIP will improve in predicting surface (total) energies while taking advantage of what it has already learned from the MPtrj dataset. We fine-tune the MACE and M3GNet models, given the availability of the corresponding implementation. We do not fine-tune the CHGNet model since the dataset does not contain atomic magnetic moments. We fine-tune the MACE model for another 200 epochs using the training set from ref. [28]. The fine-tuned model is called MACE [FT], as shown in the last line of Fig. 6 and Table 1. We include in our comparison MACE and M3GNet models trained from scratch with the same data used for fine-tuning, so-called MACE/M3GNet [Ong21]. A comparison between [FT] and [Ong21] may reveal how the model can use its previous knowledge to improve its prediction versus learning from the data from scratch.
After fine-tuning, we observe that MACE prediction of bulk total energies largely improved for most systems. It is noticeable how fine-tuning should be performed with care. As discussed before, both MTP and NequIP did not show great performance over bulk structures composed of germanium (Ge). After fine-tuning, MACE [FT] decreased its performance on germanium (Ge) bulk structures.The performance on surface total energies improved for molybdenum (Mo), while for the other elements it mostly decreased. For Li, Mo, and Si, we observe significant improvement over surface energies. Overall, the fine-tuned universal model represents an important improvement: for the selected systems, the fine-tuned MACE model is of similar performance to specialized MLIPs.
On comparison between MACE [FT] and MACE [Ong21], for Ni, Cu, and Li, their accuracy is very similar, while for Mo and Si, MACE [FT] shows significantly more accurate prediction, with the opposite for Ge. A closer look at Ge reveals that performance over bulk structures is of a similar error, however, the FT model outperforms MACE [Ong21] on surface total energies, while for surface energies, MACE [Ong21] outperforms MACE [FT].
The comparison between M3GNet and M3GNet [FT] shows smaller differences, while errors on the bulk total energies increase, there are improvements on both surface total energies and surface energies, although smaller than 20% on RMSE. On the other hand, when comparing M3GNet with M3GNet [Ong21], the benefits of the universal pretraining strategy are clearer for this model, whereas the from-scratch model shows poor performance.
II.4.3 Surface relative energy order
In addition to the RMSE for the surfaces, we investigate if MLIPs, especially UIPs, are able to accurately reproduce the relative energy order of surfaces with different orientations. This is particularly important when using UIPs for an initial screening of low-energy surfaces and structures, as well as to reproduce fundamental properties of materials, such as the Wulff shape.
Figure 7 shows the 5 lowest energy surfaces found by each MLIP in comparison with our reference data, which are the DFT surface energies. The comparison uses only the surfaces derived from the most stable bulk structure of each element, specifically mp-23, mp-30, mp-51, mp-129, mp-149, and mp-32, for Ni, Cu, Li, Mo, Si, and Ge, respectively.
Across these six elements, MACE and CHGNet tend to reproduce well the first two surfaces for Ni and Cu. All three UIPs fail to reproduce the order of the lowest surface energy for Li, Mo, and Ge. M3GNet is able to correctly predict the lowest surface energy for Si, while MACE and CHGNet fail. In fine-tuned models, MACE [FT] reproduces the lowest energy surface for Si and Ge, and M3GNet can correctly predict the lowest surface energy for Ni, Cu, and Mo. Specialized models trained with the [Ong21] dataset also struggle. For instance, M3GNet [Ong21], NequIP [Ong21], and MTP [Ong21] can correctly predict the lowest surface energy for Mo, while M3GNet [Ong21] and NequIP [Ong21] can do so for Si, and MACE [Ong21] only for Ge.